# The Complexity of Proper Homotopy Equivalence of Graphs

### v1

@@ 4. Extending the results to other spaces
Following the proof of 3.3, we just need to modify the constructions of $\Delta_{U}$ and $\Delta_{V}$ to show that the corresponding sets $U_{n}$ and $V_{n}$ are dense in each $\mathcal{G}_{k}$. For each step we start with any graph ${\Gamma}\in\mathcal{G}_{k}$ and construct $k$-regular graphs, called $\Delta_{U}^{k}$ and $\Delta_{V}^{k}$, with the desired property.
  1. (1)
    To construct $\Delta_{U}^{k}$: subdivide $n$ edges of ${\Gamma}$ outside of $B_{r}({\Gamma})$ and attach the following graphs to the new vertices.
    • For even $k$, attach $k-2$ self loops.
    • For odd $k$, attach the barrel cactus graphs from 4.1 and fig. 10. That is, for each vertex from the subdivision, add another vertex, connect them by $k-2$ edges, and add a self-loop to the new vertex.
  2. (2)
    To construct $\Delta_{V}^{k}$: For each pair of connected components of ${\Gamma}\setminus B_{r}({\Gamma})$ subdivide an edge in each and add $k-2$ edges between the new vertices.
Because a $k$-regular graph is also in $\mathcal{G}_{\leq k}$, and $\mathcal{G}_{<\infty}$, this concludes the proof for all spaces. ∎
When we defined the spaces $\mathcal{G}_{\star}$ we required that the graphs be infinite. We can drop this requirement and define analogous spaces with the same topology, call them $\mathcal{G}^{\prime}_{\star}$. These spaces will look the same as before but with the addition of the countably many finite graphs, which are all isolated points in the topology. The complexity of $\operatorname{\thicksim_{\text{PHE}}}$ does not change, since $\mathcal{G}_{\leq k}\subseteq\mathcal{G}^{\prime}_{\leq k}$ and the restriction of $\operatorname{\thicksim_{\text{PHE}}}$ to the set of finite graphs is smooth since rank is a Borel function from the space of finite graphs to $\mathbb{N}$. However, there is no longer a dense $G_{\delta}$ equivalence class, since each finite graph is an isolated point in the topology. When we construct surfaces from the finite graphs we get compact surfaces with a pants decomposition in the analogously defined spaces $\mathcal{S}^{\prime}_{\star}$. Because the homeomorphisms between $\mathcal{G}^{\prime}_{\star}$ and $\mathcal{S}^{\prime}_{\star}$ respect the equivalence classes of finite graphs and compact surfaces, which are determined only by rank and genus, we get all the analogous results for $\mathcal{S}_{\star}$. The conclusions of this discussion are summarized in the table below.
Space $(k\geq 3)$ Complexity class Dense $G_{\delta}$ equivalence class?
$\mathcal{G}_{k}$, $\mathcal{G}_{\leq k}$, $\mathcal{G}_{\leq\infty}$ $C_{\infty}$ Yes: Loch Ness Monster Graph
$\mathcal{G}^{\prime}_{k}$, $\mathcal{G}^{\prime}_{\leq k}$, $\mathcal{G}^{\prime}_{\leq\infty}$ $C_{\infty}$ No
$\mathcal{S}_{k}$, $\mathcal{S}_{\leq k}$, $\mathcal{S}_{\leq\infty}$ $C_{\infty}$ Yes: Loch Ness Monster Surface
$\mathcal{S}^{\prime}_{k}$, $\mathcal{S}^{\prime}_{\leq k}$, $\mathcal{S}^{\prime}_{\leq\infty}$ $C_{\infty}$ No
Table 1. Summary of results.

### v2

@@ Appendix A The end space map is Borel
We will induct on the length of $x$, starting with the observation that $\mathcal{H}_{\emptyset}=\mathcal{MG}_{\star}$. We will prove the stronger statement that $\mathcal{H}_{x}$ is Borel, and for any $I\subseteq\mathbb{N}$, the set $\mathcal{H}_{x,I}:=\{\Gamma\in\mathcal{H}_{x}:\forall i\in I,\;i\in U_{x}\}$ is Borel. We will use A.4 throughout.
For $|x|=1$, i.e., $x=n$ for some $n\in\mathbb{N}$, note that $\mathcal{H}_{n}$ is the collection of ${\Gamma}\in\mathcal{MG}_{\star}$ such that ${\Gamma}[>0]$ has at least $n+1$ connected components. This is an $F_{\sigma}$ set. Furthermore, $\mathcal{H}_{n,I}$ is the set of $\Gamma\in\mathcal{H}_{n}$ such that for some $m\in\mathbb{N}$ and for all $i\in I$, $m$ is the least vertex in the $(n+1)^{\text{th}}$ connected component of $\Gamma[>0]$ and each $i\in I$ is connected to $m$ in $\Gamma[>0]$. Again, this is Borel by A.4.
Now suppose $\mathcal{H}_{x}$ is Borel with $|x|=k\geq 1$, and let $n\in\mathbb{N}$. Suppose also that for any $I\subseteq\mathbb{N}$, $\mathcal{H}_{x,I}:=\{\Gamma\in\mathcal{H}_{x}:\forall i\in I,\;i\in U_{x}\}$ is Borel. $\mathcal{H}_{x^{\smallfrown}n}$ is the set of $\Gamma\in\mathcal{H}_{x}$ such that $U_{x}$ splits into at least $n+1$ connected components when we remove vertex k1k-1. That is, ${\Gamma}\in\mathcal{H}_{x^{\smallfrown}n}$ if and only if ${\Gamma}\in\mathcal{H}_{x}$ and there is an $n+1$-tuple of vertices $I=(x_{1},...,x_{n}+1)$ such that each $x_{i}\in U_{x}$, and the $x_{i}$ are pairwise disconnected in $\Gamma[>k-1]$. Thus, $\mathcal{H}_{x\smallfrown n}$ is Borel.
Finally, we need to show that for any $I\subseteq\mathbb{N}$, $\mathcal{H}_{x\smallfrown n,I}$ is Borel. Let $j$ be the smallest vertex in $I$. ${\Gamma}\in\mathcal{H}_{x\smallfrown n,I}$ if and only if
  • ${\Gamma}\in\mathcal{H}_{x^{\smallfrown}i}$,
  • for some $N$-tuple of vertices $J=\{x_{1},...,x_{N}\}$ with each $x_{i}<j$, ${\Gamma}\in\mathcal{H}_{x,I\cup J}$ and $n\leq N<\infty$,
  • for all $N+1$-tuples $J^{\prime}=\{y_{1},...,y_{N+1}\}$, with each $x_{i}<j$, ${\Gamma}\notin\mathcal{H}_{x,I\cup J}$,
  • for some $n$-tuple of vertices $K=\{z_{1},...,z_{n}\}\subseteq J$ with each $z_{i}<j$, the $z_{i}$ and $j$ are pairwise disconnected in $\Gamma[>k-1]$, and
  • for each $x_{i}\in J$, $x_{i}$ is connected to one of the $z_{k}$.
Therefore, $\mathcal{H}_{x^{\smallfrown}n,I}$ is Borel, as desired. That ${\Gamma}\mapsto\tau_{\Gamma}^{\ell}$ follows from the fact that checking whether $U_{x}$ has positive rank is Borel. ∎
Corollary A.6.
The end space maps $E_{\star},E_{\ell,\star}:\mathcal{MG}_{\star}\to\mathcal{K}(\mathbb{N}^{\mathbb{N}})$ are Borel.